i dont understand how to do this work....
solve for x
(4x+1)-(x+4)=18
i know that you have to remove par.
so i did that
and then my teacher told me to combine like terms..
that leaves me with 3x+5=18
then now i subtract 5 from both sides...
that leaves me with 3x=13
i dont think that i did that right
but i always get stuck
on certain parts
can you please help me to try to understand how to do this?
any help would be appreciated

Answer by ntnk(50)

(4x + 1) - (x + 4) = 18
Set this equation up vertically it might be easier:
4x + 1
-1x - 4
-------
3x + -3 ... REMEMBER TO SUBTRACT THE 4 FROM THE ONE, NOT ADD THEM
3x - 3 = 18
Add 3 to each side:
3x - 3 + 3 = 18 + 3
3x = 21
Divide both sides by 3:
3x/3 = 21/3
x+=+7
Check: (4(7) + 1) - (7+4) = 18
(28 +1) - (11) = 18
29 - 11 = 18
18 = 18


Solved By: NtNk